A marketing company surveys 3,200 people. They surveyed three times as many females as males. How many males and females did they survey?
Number of males surveyed =

Number of females =

Answer by ikleyn(52781)

.

x = number of males

3x = number of females.


Equation

    x + 3x = 3200

      4x   = 3200

       x   = 3200/4 = 800.


ANSWER.  800 males and 3*800 = 2400 females.

Solved.

As simple as a cucumber.